# Large Language Model for Table Processing: A Survey

### v2

@@ Abstract
Tables, typically two-dimensional and structured to store large amounts of data, are essential in daily activities like database queries, spreadsheet manipulations, web table question answering, and image table information extraction. Automating these table-centric tasks with Large Language Models (LLMs) or Visual Language Models (VLMs) offers significant public benefits, garnering interest from academia and industry. This survey provides a comprehensive overview of table-related tasks, examining both user scenarios and technical aspects. It covers traditional tasks like table question answering as well as emerging fields such as spreadsheet manipulation and table data analysis. We summarize the training techniques for LLMs and VLMs tailored for table processing. Additionally, we discuss prompt engineering, particularly the use of LLM-powered agents, for various table-related tasks. Finally, we highlight several challenges, including processing implicit user intentions and extracting information from various table sources.
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated their effectiveness and versatility across diverse tasks, leading to significant advancements in natural language processing [5]. This success has spurred researchers to investigate the application of LLMs to table-related tasks. However, the structure of tables differs from the plain text typically used during LLM pre-training. Unlike natural language corpora [6], tables possess unique characteristics.
@@
  • Structured Data: Tables are inherently structured, composed of rows and columns, each with its own schema that outlines the data’s semantics and their interrelations. Humans can effortlessly interpret tables both vertically and horizontally, but LLMs, primarily trained with sequential text data, struggle with understanding the multidimensional aspects of tables.
@@
  • Complex Reasoning: Tasks in table processing often require numerical operations (like comparisons or aggregations), data preparation (such as column type annotation and missing value detection), and more sophisticated analyses (including feature engineering and visualization). These tasks demand intricate reasoning, the ability to decompose problems into multiple steps, and logical operations, thereby posing significant challenges to machine intelligence.
@@
  • Utilizing External Tools: In real-world scenarios, humans often depend on specialized tools such as Microsoft Excel, Python, or SQL for interacting with tables. For effective table processing, LLMs need to be adept at integrating and using these external tools.
@@
The unique challenges presented by table processing tasks emphasize the need to tailor LLMs for these specific purposes. Early research, such as TaBERT [7], TaPas [8], TURL [9], and TaPEx [10], adhere to the paradigm of pre-training or fine-tuning neural language models for tables. These methods adapt model architectures, including position embeddings, attention mechanisms, and learning objectives for pretraining tasks. While these approaches yield good results, they are largely confined to specific table tasks like table question answering (table QA) and fact verification. Additionally, the BERT or BART models they utilize are not sufficiently large or versatile to handle a broader range of table tasks. Latest LLM-based approaches tackle table tasks in two primary ways: (1) curating table datasets and pre-train or fine-tune a table model [11, 12, 13]; (2) prompting an LLM or building an LLM-powered agent by utilizing the LLM’s strong reasoning ability to understand table data [14, 15, 16, 17]. These newer methods leverage LLM-specific technologies, such as instruction-tuning [18], in-context learning [19], chain-of-thought reasoning [20], and autonomous agents [21], showcasing a more versatile and comprehensive approach to table processing.
[…]
@@
Comparison with Related Surveys. Earlier surveys, such as those by Dong et al. [1] and Badaro et al. [2], primarily concentrate on pre-training or fine-tuning techniques using smaller models like BERT [8, 7] or BART [10]. However, they do not address methods based on LLMs, particularly those involving prompting strategies and agent-based approaches. Additionally, some surveys are confined to limited table tasks. For instance, Jin et al. [22] focus solely on table QA. Zhang et al. [4] focus on table reasoning, overlooking tasks such as spreadsheet manipulation. Fang et al. [3] review research on table data prediction, generation, and understanding; however, the discussion on table processing lacks depth. Qin et al. [23] concentrate on natural language to SQL (NL2SQL), overlooking spreadsheet manipulation and data analysis tasks.
@@ 2.1 Table Definition
  • Web Table (WT): The Web hosts a multitude of tables, created for diverse purposes and containing valuable information [26, 27, 24]. These tables exist in various formats, from HTML to markdown, JSON, and XML. As web tables are embedded in web pages, they have much contextual information, such as the page’s title, the surrounding text, and so on. Web tables can also have various styles and embed URLs. A well-known web table example is the table on a Wikipedia page shown in Fig. 1 (b). Web tables contain substantial factual knowledge, and the community has been working to extract and structure these tables. For example, the WDC (Web Data Commons) Web Table Corpus project converts billions of web pages from the Common Crawl corpus into structured tables [24], and Bhagavatula et al. [27] extract millions of tables from Wikipedia. The extraction process involves tasks like entity linking or relation extraction, and the extracted tables can be utilized for question answering.
@@ 2.3 Table Tasks
Table 1: Table tasks, input types, descriptions (related work), and representative datasets. In the Input Type column, the abbreviations are as follows: WT for Web Table, SS for Spreadsheet, DB for Database, and DOC for Document.
@@
Table 1 summarizes table tasks that can be automated by LLMs.
Table QA and fact verification are the most traditional table tasks, which extract knowledge from tables to answer natural language (NL) questions. Table-to-text produces an NL text based on table data. Data cleaning identifies and corrects errors in table data. Column/Row/Cell population generates possible column/row/cell for a table. Entity linking disambiguates specific entities mentioned, while column type annotation categorizes columns with types from knowledge bases. These two tasks often utilize external knowledge bases. Spreadsheet systems are originally designed for human users. Spreadsheet manipulation is a task that leverages AI to modify spreadsheets automatically, where AI accesses spreadsheet systems’ APIs or formulas. NL2SQL translates NL questions into SQL queries and can improve the efficiency of data analysts when writing SQL queries. This task has been extensively studied for years, and LLMs enhance accuracy in this field. Data analysis consists of feature engineering, machine learning, etc. Table detection identifies tables within documents, while table extraction converts them into table-native formats such as markdown, HTML, or spreadsheet. The tasks mentioned above can broadly be categorized into table-related, spreadsheet-related, database-related, and document-related tasks. These tasks require AI models to directly understand table contents, write code to manipulate spreadsheets, write SQL to access databases, or extract table data form documents.
@@ 2.4 Data Lifecycle and End-users’ Perspective
Researchers often concentrate on designing new methods to improve performance on benchmarks. However, end-users are primarily interested in how table-related AI systems can boost their productivity rather than benchmark results. To meet end-users’ needs, the industry focuses on developing products and tools for them.
@@ 3 Table Data Representation
Modern neural networks need to encode text or images into numerical representations or embeddings, and then perform computations on these embeddings. Consequently, we must format table data accordingly and feed it into language or visual language models. This section focuses on the serialization of text tables and the processing of document tables.
@@ 3.2 Visual and Layout Representation
Visual Embedding Visual embedding is the combination of image, position, and segment embeddings. To handle table images in documents, AI systems must encode images into features. For example, LayoutLM and TableVLM [44] utilize
ResNet [56] to convert images into visual embeddings, whereas Table-LLaVA [13] employs a Vision Transformer (ViT) [57].
@@ 4 Table Training
In this section, we explore the training techniques of large models for table tasks. There are primarily two types of LLMs: large language models (LLMs) that accept only text input and visual language models (VLMs) that can process visual inputs. Given the differences in inputs, model architecture, and training techniques between these two types, we summarize table training techniques in Fig. 3 and will discuss each category individually. We begin by reviewing the literature before the LLM era, i.e., models with parameter sizes smaller than 1 billion. We will then delve into the details of LLMs and VLMs.
@@ 4.2 Table LLM Training
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=3em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.5pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5em,font=,, where level=2text width=18em,font=,, where level=3text width=10em,font=,, [ Table Training, ver [ LLM,fill=blue!10 [ Continue Pre-training,fill=blue!10 [ CodeS [38], leaf, text width=5em ] ] [ Instruction Tuning, fill=blue!10 [ Table Tuning, fill=blue!10 [ TableLlama [11], Table-GPT [12], leaf, text width=12.5em ] ] [ Code Tuning, fill=blue!10 [ Magicoder [61], Lemur [42], DAAgent [17], leaf, text width=16em ] ] [ Hybrid of Table and Code, fill=blue!10 [ TableLLM [62], StructLM [63], leaf, text width=12.5em ] ] ] ] [ VLM,fill=blue!10 [ Encoder Pre-training + Fine-tuning,fill=blue!10 [ Table-LLaVA [13], TableVLM [44], TabPedia [46], leaf, text width=19em ] ] ] ]
@@ 4.2.1 What’s New in Table LLM Training
  • Table tuning focuses on LLMs’ understanding of tables so that LLMs can handle various table tasks such as table QA, table-to-text, entity linking, etc. This type of research utilizes general-purpose foundation
    LLMs (e.g., Llama) and a substantial volume of table-related data for instruction tuning. Table tuning examples include TableLlama[11] and Table-GPT[12].
@@
Continue Pre-training takes an existing trained model, feeds new data, and adapts it for a specific task [64].
@@ 4.3 Table VLM Training
Pre-training + Fine-tuning Most LLMs are designed with a decoder-only architecture. Similar to the network architecture of visual language models from the pre-LLM era, VLMs typically employ an encoder-decoder architecture, where a visual encoder converts visual data into embeddings while the decoder generates texts. The encoder may utilize architectures such as ResNet or ViT, while the decoder typically consists of a pre-trained LLM. Compared to the decoder, which has been thoroughly trained, the encoder has not yet mastered much visual information about tables. It is necessary to align the visual cues with the textual information. Therefore, the training process for a table VLM is usually divided into two phases: 1) Pre-training the visual encoder while freezing the parameters of the LLM decoder and 2) Fine-tuning or instruction tuning the entire model. Both phases require a substantial amount of high-quality training data.
4.4 Summary and Discussion
Training-based methods offer great control, allowing enterprises to conduct private training and deployment without data leakage to third-party model service providers.
Cost and Accuracy Table training on LLMs or VLMs achieves decent results and shows more generalization ability, but two issues can’t be ignored: cost and accuracy. The expense of pre-training or fine-tuning a large model is substantial. Even fine-tuning a 7B model requires eight 80GB GPUs, a cost not all institutions can afford. Moreover, the accuracy of instruction tuning LLMs is not always optimal compared with fine tuning smaller pre-trained language models, or prompting close-sourced LLMs. Li et al. [66] perform a systematic evaluation with various methods on NL2SQL task, comparing LLM-based and pre-LLM era solutions. Results show that on different metrics, different domains, there is no clear winner between LLM and pre-LLM solution. Another example is that the 7B TableLlama can’t outperform task-specific fine-tuning models on specific tasks. On data analysis tasks, instruction tuned models like Lemur and DAAgent cannot surpass GPT-4. Table-GPT continues to train on GPT-3.5 and achieves better results on all table-related tasks than GPT-3.5 and ChatGPT. However, the cost of training is prohibitively high for ordinary enterprise users who wish to deploy privately. Regarding training data, the cost of manual annotation is also high. Although synthesis is a cheap choice, the data quality is another concern. A more prevalent approach is using GPT as a teacher model for data distillation. Evidently, GPT is the performance ceiling.

### v3

@@ Abstract
Tables, typically two-dimensional and structured to store large amounts of data, are essential in daily activities like database queries, spreadsheet manipulations, web table question answering, and image table information extraction. Automating these table-centric tasks with Large Language Models (LLMs) or Visual Language Models (VLMs) offers significant public benefits, garnering interest from academia and industry. This survey provides a comprehensive overview of table-related tasks, examining both user scenarios and technical aspects. It covers traditional tasks like table question answering as well as emerging fields such as spreadsheet manipulation and table data analysis. We summarize the training techniques for LLMs and VLMs tailored for table processing. Additionally, we discuss prompt engineering, particularly the use of LLM-powered agents, for various table-related tasks. Finally, we highlight several challenges, including diverse user input when serving and slow thinking using chain-of-thought.
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated their effectiveness and versatility across diverse tasks, leading to significant advancements in natural language processing [5]. This success has spurred researchers to investigate the application of LLMs to table-related tasks. However, the structure of tables differs from the plain text [6] typically used during LLM pre-training.
@@
  • Structured Data Tables are inherently structured, composed of rows and columns, each with its own schema that outlines the data’s semantics and their interrelations. Humans can effortlessly interpret tables both vertically and horizontally, but LLMs, primarily trained with sequential text data, struggle with understanding the multidimensional aspects of tables.
@@
  • Complex Reasoning Tasks in table processing often require numerical operations (like comparisons or aggregations), data preparation (such as column type annotation and missing value detection), and more sophisticated analyses (including feature engineering and visualization). These tasks demand intricate reasoning, the ability to decompose problems into multiple steps, and logical operations, thereby posing significant challenges to machine intelligence.
@@
  • Utilizing External Tools In real-world scenarios, humans often depend on specialized tools such as Microsoft Excel, Python, or SQL for interacting with tables. For effective table processing, LLMs need to be adept at integrating and using these external tools.
Although many text-related tasks, such as those in STEM (Science, Technology, Engineering, and Mathematics) fields, require complex reasoning and external tools, table processing tasks are different due to the structural nature of tables and the user intent of querying knowledge from tables. For instance, LLMs need to understand table schemas, locate data within two-dimensional tables, and execute SQL queries to retrieve data. The unique challenges presented by table processing tasks emphasize the need to tailor LLMs for these specific purposes. Early research, such as TaBERT [7],
@@
TaPas [8], TURL [9], and TaPEx [10], adhere to the paradigm of pre-training or fine-tuning neural language models for tables. These methods adapt model architectures, including position embeddings, attention mechanisms, and learning objectives for pretraining tasks. While these approaches yield good results, they are largely confined to specific table tasks like table question answering (table QA) and fact verification. Additionally, the BERT or BART models they utilize are not sufficiently large or versatile to handle a broader range of table tasks. Latest LLM-based approaches tackle table tasks in two primary ways: (1) curating table datasets and pre-train or fine-tune a table model [11, 12, 13]; (2) prompting an LLM or building an LLM-powered agent by utilizing the LLM’s strong reasoning ability to understand table data [14, 15, 16, 17]. These newer methods leverage LLM-specific technologies, such as instruction-tuning [18], in-context learning [19], chain-of-thought reasoning [20], and autonomous agents [21], showcasing a more versatile and comprehensive approach to table processing.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=13em,font=,, where level=2text width=9em,font=,, where level=3text width=28em,font=,, where level=4text width=18em,font=,, where level=5text width=18em,font=,, [ LLM for Table Processing, ver [ Table Types and Table Tasks2), fill=blue!10 [ Table Definition2.1), fill=blue!10 [ Spreadsheet, Web Table, Database, Document, fill=blue!10 ] ] [ Table Tasks2.3), fill=blue!10 [ Table QA, Spreadsheet manipulation, NL2SQL, Table extraction, etc., fill=blue!10 ] ] ] [ Table Data Representation3), fill=orange!10 [ Text3.1), fill=orange!10 ] [ Visual and Layout3.2), fill=orange!10 ] ] [ Table Training4), fill=green!10 [ LLM4.2), fill=green!10 [ Continue Pre-training, Instruction Tuning, etc., fill=green!10 ] ] [ VLM4.3), fill=green!10 [ Encoder Pretraining + Fine-tuning, fill=green!10 ] ] ] [ Table Prompting5), fill=black!10 [ Planning5.2), fill=black!10 [ Complex Task Decomposition, Reflection & Revision, etc., fill=black!10 ] ] [ Action5.3), fill=black!10 ] ] ]
Figure 1: Taxonomy of LLMs for table processing.
Taxonomy The goal of this survey is to offer a comprehensive review of technological advancements in LLM for table processing and to summarize current research directions. As depicted in Figure 1, we have categorized the literature into a taxonomy of four key categories: table types and table tasks, table data representation, table training, and table prompting. These four categories cover distinct and interrelated research topics, offering a systematic and comprehensive review of LLM for table processing research.
Contribution The main contribution of this survey is its extensive coverage of a wide range of table tasks, including recently proposed spreadsheet manipulation and data analysis. We discuss table tasks not only from a technical perspective but also from the table data lifecycle and from the end-user’s viewpoint. We categorize methods based on the latest paradigms in LLM usage, focusing on instruction-tuning, data synthesis, chain-of-thought, ReAct, and LLM-powered agent approaches. We compile recent datasets, benchmarks, and training corpora. We collect resources such as papers, code, and datasets, which can be accessed at our website [22].
@@
Comparison with Related Surveys Earlier surveys, such as those by Dong et al. [1] and Badaro et al. [2], primarily concentrate on pre-training or fine-tuning techniques using smaller models like BERT [8, 7] or BART [10]. However, they do not address methods based on LLMs, particularly those involving prompting strategies and agent-based approaches. Additionally, some surveys are confined to limited table tasks. For instance, Jin et al. [23] focus solely on table QA. Zhang et al. [4] focus on table reasoning, overlooking tasks such as spreadsheet manipulation. Fang et al. [3] review research on table data prediction, generation, and understanding; however, the discussion on table processing lacks depth. Qin et al. [24] and Hong et al. [25] concentrate on natural language to SQL (NL2SQL), overlooking spreadsheet manipulation and data analysis tasks.
@@ 2.1 Table Definition
  • Web Table (WT): The Web hosts a multitude of tables, created for diverse purposes and containing valuable information [28, 29, 26]. These tables exist in various formats, from HTML to markdown, JSON, and XML. As web tables are embedded in web pages, they have much contextual information, such as the page’s title, the surrounding text, and so on. Web tables can also have various styles and embed URLs. A well-known web table example is the table on a
    Wikipedia page shown in Fig. 2 (b). Web tables contain substantial factual knowledge, and the community has been working to extract and structure these tables. For example, the WDC (Web Data Commons) Web Table Corpus project converts billions of web pages from the Common Crawl corpus into structured tables [26], and Bhagavatula et al. [29] extract millions of tables from Wikipedia. The extraction process involves tasks like entity linking or relation extraction, and the extracted tables can be utilized for question answering.
@@ 2.3 Table Tasks
Table 1: Table tasks, input types, descriptions (related work), and representative datasets. In the Table Type column, the abbreviations are as follows: WT for Web Table, SS for Spreadsheet, DB for Database, and DOC for Document.
@@
Table 1 summarizes table tasks that can be automated by LLMs. We also list the description of the table task, along with one or two related works, the types of tables addressed, and the datasets used.
Table QA and fact verification are the most traditional table tasks, which extract knowledge from tables to answer natural language (NL) questions. Table-to-text produces an NL text based on table data. Data cleaning identifies and corrects errors in table data. Column/Row/Cell population generates possible column/row/cell for a table. Entity linking disambiguates specific entities mentioned, while column type annotation categorizes columns with types from knowledge bases. These two tasks often utilize external knowledge bases. Spreadsheet systems are originally designed for human users. Spreadsheet manipulation is a task that leverages AI to modify spreadsheets automatically, where AI accesses spreadsheet systems’ APIs or formulas. NL2SQL translates NL questions into SQL queries and can improve the efficiency of data analysts when writing SQL queries. This task has been extensively studied for years, and LLMs enhance accuracy in this field. Data analysis consists of feature engineering, machine learning, etc. Table detection identifies tables within documents, while table extraction converts them into table-native formats such as markdown, HTML, or spreadsheet. The tasks mentioned above can broadly be categorized into table-related, spreadsheet-related,
database-related, and document-related tasks. These tasks require AI models to directly understand table contents, write code to manipulate spreadsheets, write SQL to access databases, or extract table data from documents.
@@ 2.4 Data Lifecycle and End-users’ Perspective
Researchers often concentrate on designing new
methods to improve performance on benchmarks. However, end-users are primarily interested in how table-related AI systems can boost their productivity rather than benchmark results. To meet end-users’ needs, the industry focuses on developing products and tools for them.
2.4.1 Data Lifecycle
@@
Figure 4: Three types of human interaction with table AI system: Beside, Inside, and Outside.
2.4.2 End-users’ Experience
As shown in Figure 4, from the perspective of product design and usage, the interaction between end-users and AI systems can be categorized into three types: Beside, Inside, and Outside. “Beside” refers to adding a copilot beside the application. “Inside” indicates that AI is the core component of the product. “Outside” implies that the AI system orchestrates across different applications and tasks. No matter the type, users tend to interact directly with AI through NL, and the quality of the user experience is strongly linked to the underlying table AI system [51].
@@ 3 Table Data Representation
When processing tables and table tasks, AI systems must first convert tables into machine-readable representations. Modern neural networks, which are widely adopted in AI systems, need to encode text or images into numerical representations or embeddings, and then perform computations on these embeddings. Consequently, we must format table data accordingly and feed it into language or visual language models. This section focuses on the serialization of text tables and the processing of document tables.
@@ 3.2 Visual and Layout Representation
Visual Embedding Visual embedding is the combination of image, position, and segment embeddings. To handle table images in documents, AI systems must encode images into features. For example, LayoutLM and TableVLM [46] utilize ResNet [59] to convert images into visual embeddings, whereas Table-LLaVA [13] employs a Vision Transformer (ViT) [60].
@@ 4 Table Training
In this section, we explore the training techniques of large models for table tasks. There are primarily two types of LLMs: large language models (LLMs) that accept only text input and visual language models (VLMs) that can process visual inputs. Given the differences in inputs, model architecture, and training techniques between these two types, we summarize table training techniques in Fig. 5 and will discuss each category individually. We begin by reviewing the literature on small language models (SLMs) prior to the LLM era, where these studies train models with fewer than a billion parameters. We will then delve into the details of LLMs and VLMs for tables.
@@ 4.2 Table LLM Training
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=3em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.5pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=3em,font=,, where level=2text width=13.5em,font=,, where level=3text width=10em,font=,, [ Table Training, ver [ LLM,fill=blue!10 [ Continue Pre-training,fill=blue!10 [ CodeS [40], leaf, text width=5em ] ] [ Instruction Tuning, fill=blue!10 [ Table Tuning, fill=blue!10 [ TableLlama [11], Table-GPT [12], leaf, text width=12.5em ] ] [ Code Tuning, fill=blue!10 [ Magicoder [64], Lemur [44],
DAAgent [17], SENSE [65],
FinSQL [66], leaf, text width=12.5em ] ] [ Hybrid of Table and Code, fill=blue!10 [ TableLLM [67], StructLM [68], leaf, text width=12.5em ] ] ] LLM,fill=blue!10 [ SLM fine-tuning + LLM,fill=blue!10 [ ZeroNL2SQL [69], leaf, text width=8em ] ] ] [ VLM,fill=blue!10 [ Encoder Pre-training + Fine-tuning,fill=blue!10 [ Table-LLaVA [13], TableVLM [46], TabPedia [48], PixT3 [70], leaf, text width=24em ] ] ] ]
@@ 4.2.1 What’s New in Table LLM Training
  • Table Tuning focuses on LLMs’ understanding of tables so that LLMs can handle various table tasks such as table QA, table-to-text, entity linking, etc. This type of research utilizes general-purpose foundation LLMs
    (e.g., Llama) and a substantial volume of table-related data for instruction tuning. Table tuning examples include TableLlama[11] and
    Table-GPT[12].
@@
Continue Pre-training takes an existing trained model, feeds new data, and adapts it for a specific task [71]. SLMs + LLMs is a paradigm to fine-tune an SLM to guide the LLM towards desired outputs where two types of models complement each other [72].
@@ 4.2.3 Code Tuing
SENSE [65] is an NL2SQL model that utilizes two types of synthetic training data: “strong data” and “weak data.” “Strong data” is distilled from powerful LLMs that provide more reliable responses, while “weak data” is produced by inferior models that may result in errors during SQL execution. The strong and weak data are then trained with Direct Preference Optimization (DPO) [74], enabling the SENSE model to learn from both correct and incorrect samples. FinSQL is also an NL2SQL solution. It creates a dataset specialized for the financial sector and employs a powerful LLM to augment the data, followed by parameter-efficient fine-tuning (PEFT) [75, 76] of a LLM.
@@ 4.2.5 Continue Pre-training
4.2.6 SLMs + LLMs
SLMs are easier to fine-tune to understand table schemas, whereas LLMs exhibit strong reasoning capabilities but may encounter “hallucination” [79] problems owing to the lack of domain knowledge. ZeroNL2SQL [69] combines the strengths of both SLMs and LLMs to mitigate their respective weaknesses. It fine-tunes an Encoder-Decoder SLM responsible for generating SQL sketch candidates, and it employs an LLM to fill in missing parts in the SQL sketch, correct errors in the SQL query, and generate the final query.
@@ 4.3 Table VLM Training
Pre-training + Fine-tuning Most LLMs are designed with a decoder-only architecture. Similar to the network architecture of visual language models from the pre-LLM era, VLMs typically employ an encoder-decoder architecture, where a visual encoder converts visual data into embeddings while the decoder generates texts. The encoder may utilize architectures such as ResNet or ViT, while the decoder typically consists of a pre-trained LLM. Compared to the decoder, which has been thoroughly trained, the encoder has not yet mastered much visual information about tables. It is necessary to align the visual cues with the textual information. Therefore, the training process for a table VLM is usually divided into two phases: 1) Pre-training the visual encoder while freezing the parameters of the LLM decoder and 2) Fine-tuning or instruction tuning the entire model. Both phases require a substantial amount of high-quality training data. PixT3 [70] is a multimodal table-to-text model that takes table-to-text tasks as table visual recognition tasks and generates texts, removing the need to process tables in text formats.